OpenStudy (anonymous):

Diana took her team to breakfast to celebrate Sonia’s birthday. Brian’s breakfast cost $9.50, Steve’s breakfast cost $7.00. Both Ken and Judy’s breakfast cost $6.50 each. Diana’s breakfast cost $7.50. Sonia’s breakfast, being that it was her birthday, was free. If the sales tax in Palm Beach County is 6%, how much was the total bill, including the sales tax?

OpenStudy (anonymous):

(9.5+7.0+6.5 + 6.5 +7.5)×1.06= You would total the cost for each of the diners then you would add the sales tax. You can multiply the sum by .06 and then add it back to the sum or to make it easy just multiply by 1.06. The answer in this case would be $39.22

OpenStudy (anonymous):

9.50 + 7.00 + 6.50 + 6.50 + 7.50 = 37.00 (cost of food without sales tax) 6% = .06 (sales tax turned to a decimal) .06 x 37.00 = 2.22 (amount of sales tax) 2.22 + 37.00 = 39.22 ( total bill including sales tax)